OpenStudy (anonymous):

51a+2b=84c How to find smallest solutions for a,b,c

OpenStudy (anonymous):

@ParthKohli any ideas?

Parth (parthkohli):

Diophantine equations. Gotta hate them.

OpenStudy (anonymous):

I dont think the question is complete

Parth (parthkohli):

Well, do you need positive integer solutions?

OpenStudy (anonymous):

Yep. Smallest positive

Parth (parthkohli):

Oh yeah, how do you define smallest \(a,b,c\)?

OpenStudy (anonymous):

a b c are integers. no common factors

OpenStudy (anonymous):

did u mean a+b+c to be smallest

OpenStudy (anonymous):

that also works

Parth (parthkohli):

Oh, you need coprime a,b,c.

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Can someone explain me what the real question is?

OpenStudy (anonymous):

what is the smallest amounts of 51 and 2 required to make a multiple of 84

Parth (parthkohli):

lol, I have \(a = b = c= 0\). ;-)

OpenStudy (anonymous):

perhaps I should have mentioned positive :D

OpenStudy (anonymous):

51*2 + 2*33=84*2

OpenStudy (anonymous):

Oh wait...I did

Parth (parthkohli):

@sauravshakya Nailed it.

OpenStudy (anonymous):

Note: a must be even

OpenStudy (callisto):

smallest amount of 52 // smallest amount of 2 // small amount of 52 and 2?

Parth (parthkohli):

\[51a + 2b \equiv 0\pmod{84}\]So yup, \(a\) can't be odd.

OpenStudy (callisto):

That means smallest a / smallest b/ smallest c / smallest a+b ?

OpenStudy (anonymous):

8(51)+6(2) = 5(84)

OpenStudy (anonymous):

So u want to find minimum value of a+b+c???

OpenStudy (anonymous):

yes

Parth (parthkohli):

But they aren't coprime. :-\

OpenStudy (anonymous):

Oh whoops, then I didn't mean coprime As in like a,b,c don't have any common factors across the three of them

OpenStudy (anonymous):

It's all good then guys :D